Find an equation for the line passing through the point and parallel to the given line: (1,2) Y= 3x-2
Answer by stanbon(75887)

Find an equation for the line passing through the point and parallel to the given line: (1,2) Y= 3x-2
---------------------------
The slope of the given line is 3.
Any line parallel to it must have a slope of 3.
Find the intercept:
since y = mx+b, 2 = 3*1+b
b = 2-3 = -1
----------------------
Equation:
y = 3x-1
==================
Cheers,
Stan H.